Last night Rana noticed that the overflows on the ITM and ETM coils were a crazy huge number. Today I rebooted c1dcuepics, c1iovme, c1sosvme, c1susvme1 and c1susvme2 (in that order). Rob helped me burt restore losepics and iscepics, which needs to be done whenever you reboot the epics computer.
Unfortunately this didn't help the overflow problem at all. I don't know what to do about that.
Just start by re-setting them to zero. Then you have to figure out what's causing them to saturate by watching time series and looking at spectra.